Please solve the equation for x: 3/4x+2=x-1
Answer by checkley77(12844)

3/4x+2=x-1
3X/4+2=X-1
3X/4-X=-1-2
(3X-4X)/4=-3
-X/4=-3 CROSS MULTIPLY
-X=-3*4
-X=-12
X=12 ANS.
PROOF:
3*12/4+2=12-1
36/4+2=11
9+2=11
11=11